OpenStudy (anonymous):

x^5y^-2x/x^3y^5z^-3 Please help!

OpenStudy (jackellyn):

Hey @JulieeeXD , when dividing variables with exponents you're basically subtracting the exponents and keeping the variables. So \[\frac{ x^5y^{-2} }{ x^3y^5z^{-3} }\] x: 5-3=2 y: -2-5=-7 z: -3 \[x^2y^{-7}z^{-3}\]